If f(x)=5x-1 , find the following
f(a-2)
f(a-2)=

Answer by rapaljer(4671)

If f(x) = 5x-1, then to find f(a-2), you must substitute the a-2 in place of all the x's.

f(x) = 5(x) - 1
f(a-2) = 5(a-2) -1

Now, simplify the expression:
f(a-2) = 5a-10 -1
f(a-2) = 5a -11

R^2 at SCC